OpenStudy (anonymous):

Solve for x x^2+50=10x

OpenStudy (anonymous):

I got x=5+5i, 5-5i I need the answer in the form a+bi, how do I do that?

myininaya (myininaya):

x^2+50-10x=10x-10x x^2-10x+50=0 using ax^2+bx+c=0 where a=1, b=-10 c=50 \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-(-10) \pm \sqrt{(-10)^2-4(1)(50)}}{2(1)}\]

OpenStudy (anonymous):

so that is the answer then myininaya

myininaya (myininaya):

\[x=\frac{10 \pm \sqrt{100-200}}{2}=\frac{10 \pm \sqrt{-100}}{2}=\frac{10 \pm i10}{2}\] yes you are right

OpenStudy (anonymous):

which one then the top or bottom

myininaya (myininaya):

what do you mean?

OpenStudy (anonymous):

the top answer or the bottom won before u say ur right

myininaya (myininaya):

its all the same

OpenStudy (anonymous):

ok

myininaya (myininaya):

but i said you were right (10+10i)/2 is 5+5i and 10-10i)/2 is 5-5i